OpenStudy (anonymous):

Yes i do not know how to do this. I'm going out so my respond is going to be late. Can someone explain how to do this in steps not just a list of things for me to figure out. Thanks! For y = 4x + 2x2 – x3, show how to find the value for y, when x = 2. Show your work y = 2x + 2x^2+ -x^3 y3x^-1 -x^3

OpenStudy (anonymous):

y=2x+2x^2 - x^3 first replace all of the x's with 2 y=2(2) +2(2)^2 - 2^3 deal with your exponents y=2(2) +2(4) - 8 multiply y= 4+8 -8 simplify y= 4

OpenStudy (anonymous):

Shouldn't 2(2) = 4

OpenStudy (anonymous):

2(2)=4 + 2(2)^2 = 8 and -2^3 = 6

OpenStudy (anonymous):

Not getting the second 8.

OpenStudy (anonymous):

?

OpenStudy (lynfran):

\[y=4x+2x ^{2}-x ^{3} \] is this the problem?

OpenStudy (anonymous):

Yeah, didn't notice i changed the 4 into a two sh*t

OpenStudy (anonymous):

4(2)+2(2)^2 -2^3 This right?

OpenStudy (lynfran):

so to go about finding for y, we sub substitute the value of x...which is 2... so\[y=4(2)+2(2)^{2}-(2)^{3}\]

OpenStudy (lynfran):

we ditribute and get \[y=8+2(4)-8\]

OpenStudy (anonymous):

Okay thanks i think i got it.

OpenStudy (lynfran):

simplify more we get\[y=8+8-8\]..ok whats your answer?

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

4(2)+2(2)^2 -2^3 4(2) = 8, +2(2) =4,^2 =8, -2^3 =8, 8+8 -8

OpenStudy (lynfran):

ok i what you to tell me what y=? ...8+8-8 can be simplify

OpenStudy (anonymous):

y=8

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

Thanks :)

OpenStudy (lynfran):

welcome